# Double Robustness of Local Projections

### v3

@@ 1 Introduction
While this sentiment has been echoed in influential reviews (e.g., Ramey, 2016; Nakamura and Steinsson, 2018; Jordà, 2023), there so far exist essentially no theoretical results on the relative robustness of LP and VAR inference procedures to misspecification. Plagborg-Møller and Wolf (2021) and Xu (2023) show that the two estimators are in fact asymptotically equivalent—and thus equally robust to misspecification—in a general VAR($\infty$) model if the estimation lag length diverges to infinity with the sample size. However, this result does not directly speak to the empirically relevant case where researchers employ small-to-moderate lag lengths to preserve degrees of freedom. Applied researchers must therefore base their choice of inference procedure on empirically calibrated simulation studies (Kilian and Kim, 2011; Li, Plagborg-Møller, and Wolf, 2024).
@@
In this setting, we prove that the conventional LP confidence interval has correct (pointwise) asymptotic coverage even for local misspecification that is of such a large magnitude that it can be detected with probability 1 in large samples. This robustness property requires that we control for those lags of the data that are strong predictors of the outcome or impulse variables, but—crucially for applied work—the omission of lags with small-to-moderate predictive power does not threaten coverage. We argue that our result can be interpreted as a consequence of the double robustness of the LP estimator, which is analogous to the double robustness of modern partially linear regression estimators in the literature on debiased machine learning. [^1] Important contributions include Robins, Mark, and Newey (1992), Robins and Rotnitzky (1995), Robins, Rotnitzky, and van der Laan (2000), Robins and Rotnitzky (2001), Bang and Robins (2005), Ai and Chen (2007), Chernozhukov, Chetverikov, Demirer, Duflo, Hansen, Newey, and Robins (2018), and Chernozhukov, Escanciano, Ichimura, Newey, and Robins (2022).
@@
Relative to the previously cited simulation studies of LPs and VARs, we here derive analytical results on the worst-case asymptotic properties of these two inference procedures that hold for a wide range of stationary, locally misspecified VAR models. The simulations in Li, Plagborg-Møller, and Wolf (2024) suggest a stark bias-variance trade-off between LP (low bias, high variance) and moderate-lag VAR estimators (moderate bias, low variance). The reason behind the theoretical superiority of LP proved in this paper is that, if the objective is to construct confidence intervals with robust coverage for a wide range of DGPs, then even a moderate amount of VAR bias cannot be tolerated, as it causes the VAR confidence interval to be poorly centered. A concern for correct confidence interval coverage thus effectively induces a large weight on bias in the researcher’s objective function, justifying the use of LP despite its higher variance.
@@
We also build upon previous research into misspecified VAR models, uncovering novel results about the robustness of LPs and the worst-case properties of VAR procedures. Braun and Mittnik (1993) derive expressions for the probability limits of VAR estimators under global MA misspecification; however, since bias always dominates variance asymptotically in their framework, they do not characterize the properties of LP and VAR inference procedures, which is the focus of our paper. Schorfheide (2005) characterizes the asymptotic mean squared errors of iterated and direct multi-step forecasts in a reduced-form VAR model with MA terms of order $T^{-1/2}$, and González-Casasús and Schorfheide (2025) use this framework to select hyperparameters for VAR forecasts. Müller and Stock (2011) construct Bayesian forecast intervals in a locally misspecified univariate AR model. Relative to these papers, we here contribute by: (i) focusing on structural impulse responses rather than forecasting; (ii) allowing for more general rates of local misspecification, which is key to uncovering the double robustness of LP; and (iii) deriving simple analytical formulae for worst-case bias and coverage of VARs. As such, our results formalize concerns by applied practitioners about the lack of VAR robustness and sensitivity to lag length (Chari, Kehoe, and McGrattan, 2008; Nakamura and Steinsson, 2018; see also Inoue and Kilian, 2002, and Kilian and Lütkepohl, 2017, Chapters 2.6.5 and 6.2).
@@ 2.1 Model and assumptions
The assumptions that the shocks are mutually and serially independent are made to simplify the exposition; we prove formally in LABEL:app:hetero that our results on the robustness of LP and on the asymptotic bias of VAR go through for a large class of conditionally heteroskedastic shock processes. The assumptions on $H$ correspond to recursive (also known as Cholesky) identification of the shock of interest $\varepsilon_{j^{*},t}$, with a unit effect normalization $H_{j^{*},j^{*}}=1$. A special case is when the shock is directly observed, which corresponds to ordering it first (i.e., $j^{*}=1$). LABEL:app:proxy shows that our results extend also to identification via external instruments or proxies (Stock and Watson, 2018). Absolute summability of $\alpha(L)$ is a weak regularity condition ensuring the vector MA($\infty$) process $\alpha(L)\varepsilon_{t}$ is well-defined (Brockwell and Davis, 1991, Proposition 3.1.1). Finally, the assumption that $\zeta>0$ restricts attention to local misspecification, as discussed earlier.
@@
Our local-to-SVAR model (2.1) with MA misspecification covers several empirically relevant types of model misspecification. While essentially all modern discrete-time, linearized macro models have VARMA representations, they usually cannot be represented exactly as finite-order VAR models (e.g., Kilian and Lütkepohl, 2017, Chapter 6.2). Even if the true DGP were a finite-order VAR, dynamic misspecification of the estimation model can give rise to MA terms, for example due to under-specification of the lag length or failing to control for some of the variables in the true system. Relatedly, MA terms may appear because of a failure of invertibility of the shocks (Alessi, Barigozzi, and Capasso, 2011). VARMA representations can also arise from temporal or cross-sectional aggregation of finite-order VAR models, including contamination by classical measurement error (Granger and Morris, 1976; Lütkepohl, 1984). In all of these cases, if the number of lags used for estimating the VAR is chosen to be sufficiently large, then the MA remainder will be small, consistent with the spirit of our locally misspecified model (2.1).
In terms of structural shock identification, our framework accommodates both the case of a well-identified shock (or instrument/proxy, see LABEL:app:proxy) but misspecification in other parts of the model, as well as misspecification in the structural shock identification itself. Key to this generality is that we allow the $m\times m$ MA polynomial $\alpha(L)$ to be arbitrary. To see this, consider the case $j^{*}=1$, so interest centers on the dynamic causal effects of the first shock $\varepsilon_{1,t}$. If the first row of $\alpha(L)$ is zero, then $\varepsilon_{1,t}$ is well-identified as the reduced-form residual in the first equation of the VAR. If the first row of $\alpha(L)$ is non-zero, then the reduced-form residual will be contaminated by lagged shocks, thus allowing for the possibility that shock identification itself is not entirely accurate.
@@ 2.2 Estimators
Note that the two estimators coincide at the impact horizon: $\hat{\beta}_{0}=\hat{\delta}_{0}$ (see LABEL:thm:impact in LABEL:app:proof_details).
@@ 3.1 Large-sample distributions and confidence interval coverage
Here $z_{1-a/2}$ is the 1a/21-a/2 quantile of the standard normal distribution, and $\operatorname{aVar}(\hat{\beta}_{h})$ and $\operatorname{aVar}(\hat{\delta}_{h})$ are the asymptotic variances of the leading (order-$T^{-1/2}$) stochastic terms in the representations of the LP and VAR estimators in Propositions˜3.1 and 3.2; explicit formulae for the asymptotic variances are given in Corollary˜A.2 in Section˜A.3, which also implies that $\operatorname{aVar}(\hat{\beta}_{h})\geq\operatorname{aVar}(\hat{\delta}_{h})$. None of the results below would change if we replaced the asymptotic variances with the conventional consistent estimates of these (that assume correct specification, as implemented in standard econometric software packages). [^6] Under Assumption 2.1, homoskedastic standard errors suffice. For LP, Heteroskedasticity and Autocorrelation Robust inference would generally be required under the weaker assumptions in LABEL:app:hetero, though simple heteroskedasticity-robust standard errors suffice under the assumptions discussed by Montiel Olea and Plagborg-Møller (2021) and Xu (2023).
@@ 3.3 Lag length selection
LABEL:app:info_crit shows that the LP confidence interval maintains correct asymptotic coverage when the lag length $p$ is selected via the Bayesian Information Criterion (BIC), provided that $\zeta\geq 1/2$. Specifically, the BIC should be applied to an auxiliary VAR in the observed data series $\check{y}_{t}$; the selected lag length then determines the number $p$ of lags to control for in subsequent LP inference (which otherwise discards the auxiliary VAR). The resulting LP inference is robust to model selection errors that are known to cause difficulties for VAR inference (Leeb and Pötscher, 2005; Kilian and Lütkepohl, 2017, chapter 2.6.5). At a high level, the greater reliability of LP inference with data-dependent lag length is a consequence of the double robustness discussed earlier, see Belloni and Chernozhukov (2013) and Chernozhukov et al. (2018).
@@ 4.1 Worst-case bias and mean-squared error
The claim is a special case of LABEL:thm:bias_general in LABEL:app:joint. ∎
@@ 4.2 Worst-case coverage
The shape and magnitude of the least favorable misspecification is difficult to rule out generally based on economic theory. The least favorable MA polynomial $\alpha^{\dagger}(L;h,M)=\sum_{\ell=1}^{\infty}\alpha_{\ell,h,M}^{\dagger}L^{\ell}$ for VAR coverage is the same as the least favorable one for bias (i.e., the $\alpha(L)$ that achieves the maximum in Proposition˜4.1). Since $\operatorname{aBias}(\hat{\delta}_{h})$ is linear in $\alpha(L)$, the least favorable choice given the constraint $\|\alpha(L)\|\leq M$ follows from the Cauchy-Schwarz inequality (see the proof of LABEL:thm:bias_general in LABEL:app:joint):
@@ 4.4 Inference on multiple impulse responses
Since the least favorable MA polynomial derived in Section˜4.2 depends on the horizon $h$ of interest, one might hope that VARs would not be as prone to bias and thereby undercoverage if interest centers on multiple impulse responses. Unfortunately, LABEL:app:joint shows that this is not the case. There we consider inference on a vector of impulse responses for any combination of response variables $i$, shocks $j$, and horizons $h$. Generalizing Proposition˜4.1, we show that the worst-case norm of the bias is non-negligible if the VAR offers efficiency gains for any linear combination of the parameters of interest. This implies that the conventional VAR confidence interval has fragile coverage even if the target parameter is a linear combination of impulse responses (such as the integral or sum across multiple horizons, as in the fiscal multiplier applications reviewed in Ramey, 2016). The conventional Wald confidence ellipsoid centered at the VAR estimator is similarly fragile.
@@ 5.2 Empirically calibrated simulation study
The parameters of interest are the impulse responses of CPI to the observed oil shock. To be consistent with our theory, we use an “internal instruments” specification that orders the proxy first in the VAR; this differs from the “external instruments” specification used by Känzig. The sample size is $T=720$ months. We will entertain different choices of the estimation lag length $p$ for the LP and VAR estimators. We report results for both delta method and bootstrap confidence intervals. Results are based on 10,000 Monte Carlo simulations. See LABEL:app:oil for implementation details.
@@ Results.
LABEL:app:oil illustrates that the VAR under-coverage can be ameliorated by increasing the lag length beyond what is typically used in current applied practice, at the expense of higher variance, consistent with Section˜3.3. The supplement also reports that the VAR estimator achieves lower MSE than LP, suggesting that—despite the poor performance of the VAR confidence interval—the larger bias of the VAR estimator may not compromise its usefulness as a point estimator (see also Li, Plagborg-Møller, and Wolf, 2024).
Montiel Olea, Plagborg-Møller, Qian, and Wolf (2025) find that the qualitative conclusions above extend to a wide range of empirically calibrated simulation DGPs based on richly specified dynamic factor models: VAR confidence intervals fail to adequately control coverage in a sizable fraction of DGPs, while LP confidence intervals robustly maintain accurate coverage. They document that LP is not only more robust to lag length selection, but also to the choice of control variables, consistent with the theory in this paper.
@@ 6 Conclusion
    When the goal is to construct confidence intervals for impulse responses that have accurate coverage in a wide range of empirically relevant DGPs—as opposed to minimizing MSE—then the smaller bias of LPs documented in simulations by Li, Plagborg-Møller, and Wolf (2024) is more valuable than the smaller variance enjoyed by VAR estimators.
@@
Is there a way forward for VAR inference, beyond just including a large number of lags? We showed how to construct a VAR confidence interval with a bias-aware critical value that robustly controls coverage, but found that it will typically lead to wider confidence intervals than LP. Another option would be to estimate VARMA models rather than pure VARs, though this would be computationally expensive, and the bias-variance trade-off relative to LPs is unclear. In principle, VAR procedures may work better under additional restrictions on the misspecification, such as shape restrictions on the impulse response functions. [^11] Given any convex parameter space for the misspecification MA polynomial $\alpha(L)$, the worst-case bias of the VAR estimator (see Proposition 3.2) can be computed using convex programming. However, it appears that detailed application-specific restrictions would be required to generate a negligible worst-case bias, since we have shown that the least favorable misspecification in our baseline analysis cannot generally be ruled out based on economic theory. Rather than restricting the parameter space, future research could instead investigate weakening the coverage requirement, e.g., only requiring a certain coverage probability on average over a set of horizons (Armstrong, Kolesár, and Plagborg-Møller, 2022), or by changing the target for inference from the true impulse response function to a smooth projection of this function (Genovese and Wasserman, 2008). Finally, a subjectivist Bayesian VAR modeler need only worry about our negative results if their prior on potential misspecification attaches significant weight to MA processes that imply large VAR biases.

### v4

@@ 1 Introduction
While this sentiment has been echoed in influential reviews (e.g., Ramey, 2016; Nakamura and Steinsson, 2018; Jordà, 2023), there so far exist essentially no theoretical results on the relative robustness of LP and VAR inference procedures to misspecification. Plagborg-Møller and Wolf (2021) and Xu (2023) show that the two estimators are in fact asymptotically equivalent—and thus equally robust to misspecification—in a general VAR($\infty$) model if the estimation lag length diverges to infinity with the sample size. However, this result does not directly speak to the empirically relevant case where researchers employ small-to-moderate lag lengths to preserve degrees of freedom. Applied researchers must therefore base their choice of inference procedure on empirically calibrated simulation studies (Kilian and Kim, 2011; Li et al., 2024).
@@
In this setting, we prove that the conventional LP confidence interval has correct (pointwise) asymptotic coverage even for local misspecification that is of such a large magnitude that it can be detected with probability 1 in large samples. This robustness property requires that we control for those lags of the data that are strong predictors of the outcome or impulse variables, but—crucially for applied work—the omission of lags with small-to-moderate predictive power does not threaten coverage. We argue that our result can be interpreted as a consequence of the double robustness of the LP estimator, which is analogous to the double robustness of modern partially linear regression estimators in the literature on debiased machine learning. [^1] Important contributions include Robins et al. (1992), Robins and Rotnitzky (1995), Robins et al. (2000), Robins and Rotnitzky (2001), Bang and Robins (2005), Ai and Chen (2007), Chernozhukov et al. (2018), and Chernozhukov et al. (2022).
@@
Relative to the previously cited simulation studies of LPs and VARs, we here derive analytical results on the worst-case asymptotic properties of these two inference procedures that hold for a wide range of stationary, locally misspecified VAR models. The simulations in Li et al. (2024) suggest a stark bias-variance trade-off between LP (low bias, high variance) and moderate-lag VAR estimators (moderate bias, low variance). The reason behind the theoretical superiority of LP proved in this paper is that, if the objective is to construct confidence intervals with robust coverage for a wide range of DGPs, then even a moderate amount of VAR bias cannot be tolerated, as it causes the VAR confidence interval to be poorly centered. A concern for correct confidence interval coverage thus effectively induces a large weight on bias in the researcher’s objective function, justifying the use of LP despite its higher variance.
@@
We also build upon previous research into misspecified VAR models, uncovering novel results about the robustness of LPs and the worst-case properties of VAR procedures. Braun and Mittnik (1993) derive expressions for the probability limits of VAR estimators under global MA misspecification; however, since bias always dominates variance asymptotically in their framework, they do not characterize the properties of LP and VAR inference procedures, which is the focus of our paper. Schorfheide (2005) characterizes the asymptotic mean squared errors of iterated and direct multi-step forecasts in a reduced-form VAR model with MA terms of order $T^{-1/2}$, and González-Casasús and Schorfheide (2025) use this framework to select hyperparameters for VAR forecasts. Müller and Stock (2011) construct Bayesian forecast intervals in a locally misspecified univariate AR model. Relative to these papers, we here contribute by: (i) focusing on structural impulse responses rather than forecasting; (ii) allowing for more general rates of local misspecification, which is key to uncovering the double robustness of LP; and (iii) deriving simple analytical formulae for worst-case bias and coverage of VARs. As such, our results formalize concerns by applied practitioners about the lack of VAR robustness and sensitivity to lag length (Chari et al., 2008; Nakamura and Steinsson, 2018; see also Inoue and Kilian, 2002, and Kilian and Lütkepohl, 2017, Chapters 2.6.5 and 6.2).
@@ 2.1 Model and assumptions
The assumptions that the shocks are mutually and serially independent are made to simplify the exposition; we prove formally in Supplemental Appendix C.1 that our results on the robustness of LP and on the asymptotic bias of VAR go through for a large class of conditionally heteroskedastic shock processes. The assumptions on $H$ correspond to recursive (also known as Cholesky) identification of the shock of interest $\varepsilon_{j^{*},t}$, with a unit effect normalization $H_{j^{*},j^{*}}=1$. A special case is when the shock is directly observed, which corresponds to ordering it first (i.e., $j^{*}=1$). Supplemental Appendix C.2 shows that our results extend also to identification via external instruments or proxies (Stock and Watson, 2018). Absolute summability of $\alpha(L)$ is a weak regularity condition ensuring the vector MA($\infty$) process $\alpha(L)\varepsilon_{t}$ is well-defined (Brockwell and Davis, 1991, Proposition 3.1.1). Finally, the assumption that $\zeta>0$ restricts attention to local misspecification, as discussed earlier.
@@
Our local-to-SVAR model (2.1) with MA misspecification covers several empirically relevant types of model misspecification. While essentially all modern discrete-time, linearized macro models have VARMA representations, they usually cannot be represented exactly as finite-order VAR models (e.g., Kilian and Lütkepohl, 2017, Chapter 6.2). Even if the true DGP were a finite-order VAR, dynamic misspecification of the estimation model can give rise to MA terms, for example due to under-specification of the lag length or failing to control for some of the variables in the true system. Relatedly, MA terms may appear because of a failure of invertibility of the shocks (Alessi et al., 2011). VARMA representations can also arise from temporal or cross-sectional aggregation of finite-order VAR models, including contamination by classical measurement error (Granger and Morris, 1976; Lütkepohl, 1984). In all of these cases, if the number of lags used for estimating the VAR is chosen to be sufficiently large, then the MA remainder will be small, consistent with the spirit of our locally misspecified model (2.1).
In terms of structural shock identification, our framework accommodates both the case of a well-identified shock (or instrument/proxy, see Supplemental Appendix C.2) but misspecification in other parts of the model, as well as misspecification in the structural shock identification itself. Key to this generality is that we allow the $m\times m$ MA polynomial $\alpha(L)$ to be arbitrary. To see this, consider the case $j^{*}=1$, so interest centers on the dynamic causal effects of the first shock $\varepsilon_{1,t}$. If the first row of $\alpha(L)$ is zero, then $\varepsilon_{1,t}$ is well-identified as the reduced-form residual in the first equation of the VAR. If the first row of $\alpha(L)$ is non-zero, then the reduced-form residual will be contaminated by lagged shocks, thus allowing for the possibility that shock identification itself is not entirely accurate.
@@ 2.2 Estimators
Note that the two estimators coincide at the impact horizon: $\hat{\beta}_{0}=\hat{\delta}_{0}$ (see Lemma E.5 in Supplemental Appendix E).
@@ 3.1 Large-sample distributions and confidence interval coverage
Here $z_{1-a/2}$ is the 1a/21-a/2 quantile of the standard normal distribution, and $\operatorname{aVar}(\hat{\beta}_{h})$ and $\operatorname{aVar}(\hat{\delta}_{h})$ are the asymptotic variances of the leading (order-$T^{-1/2}$) stochastic terms in the representations of the LP and VAR estimators in Propositions˜3.1 and 3.2; explicit formulae for the asymptotic variances are given in Corollary˜A.2 in Section˜A.3, which also implies that $\operatorname{aVar}(\hat{\beta}_{h})\geq\operatorname{aVar}(\hat{\delta}_{h})$. None of the results below would change if we replaced the asymptotic variances with the conventional consistent estimates of these (that assume correct specification, as implemented in standard econometric software packages). [^6] Under Assumption 2.1, homoskedastic standard errors suffice. For LP, Heteroskedasticity and Autocorrelation Robust inference would generally be required under the weaker assumptions in Supplemental Appendix C.1, though simple heteroskedasticity-robust standard errors suffice under the assumptions discussed by Montiel Olea and Plagborg-Møller (2021) and Xu (2023).
@@ 3.3 Lag length selection
Supplemental Appendix C.3 shows that the LP confidence interval maintains correct asymptotic coverage when the lag length $p$ is selected via the Bayesian Information Criterion (BIC), provided that $\zeta\geq 1/2$. Specifically, the BIC should be applied to an auxiliary VAR in the observed data series $\check{y}_{t}$; the selected lag length then determines the number $p$ of lags to control for in subsequent LP inference (which otherwise discards the auxiliary VAR). The resulting LP inference is robust to model selection errors that are known to cause difficulties for VAR inference (Leeb and Pötscher, 2005; Kilian and Lütkepohl, 2017, chapter 2.6.5). At a high level, the greater reliability of LP inference with data-dependent lag length is a consequence of the double robustness discussed earlier, see Belloni and Chernozhukov (2013) and Chernozhukov et al. (2018).
@@ 4.1 Worst-case bias and mean-squared error
The claim is a special case of Proposition C.2 in Supplemental Appendix C.4. ∎
@@ 4.2 Worst-case coverage
The shape and magnitude of the least favorable misspecification is difficult to rule out generally based on economic theory. The least favorable MA polynomial $\alpha^{\dagger}(L;h,M)=\sum_{\ell=1}^{\infty}\alpha_{\ell,h,M}^{\dagger}L^{\ell}$ for VAR coverage is the same as the least favorable one for bias (i.e., the $\alpha(L)$ that achieves the maximum in Proposition˜4.1). Since $\operatorname{aBias}(\hat{\delta}_{h})$ is linear in $\alpha(L)$, the least favorable choice given the constraint $\|\alpha(L)\|\leq M$ follows from the Cauchy-Schwarz inequality (see the proof of Proposition C.2 in Supplemental Appendix C.4):
@@ 4.4 Inference on multiple impulse responses
Since the least favorable MA polynomial derived in Section˜4.2 depends on the horizon $h$ of interest, one might hope that VARs would not be as prone to bias and thereby undercoverage if interest centers on multiple impulse responses. Unfortunately, Supplemental Appendix C.4 shows that this is not the case. There we consider inference on a vector of impulse responses for any combination of response variables $i$, shocks $j$, and horizons $h$. Generalizing Proposition˜4.1, we show that the worst-case norm of the bias is non-negligible if the VAR offers efficiency gains for any linear combination of the parameters of interest. This implies that the conventional VAR confidence interval has fragile coverage even if the target parameter is a linear combination of impulse responses (such as the integral or sum across multiple horizons, as in the fiscal multiplier applications reviewed in Ramey, 2016). The conventional Wald confidence ellipsoid centered at the VAR estimator is similarly fragile.
@@ 5.2 Empirically calibrated simulation study
The parameters of interest are the impulse responses of CPI to the observed oil shock. To be consistent with our theory, we use an “internal instruments” specification that orders the proxy first in the VAR; this differs from the “external instruments” specification used by Känzig. The sample size is $T=720$ months. We will entertain different choices of the estimation lag length $p$ for the LP and VAR estimators. We report results for both delta method and bootstrap confidence intervals. Results are based on 10,000 Monte Carlo simulations. See Supplemental Appendix D for implementation details.
@@ Results.
Supplemental Appendix D illustrates that the VAR under-coverage can be ameliorated by increasing the lag length beyond what is typically used in current applied practice, at the expense of higher variance, consistent with Section˜3.3. The supplement also reports that the VAR estimator achieves lower MSE than LP, suggesting that—despite the poor performance of the VAR confidence interval—the larger bias of the VAR estimator may not compromise its usefulness as a point estimator (see also Li et al., 2024).
Montiel Olea et al. (2025) find that the qualitative conclusions above extend to a wide range of empirically calibrated simulation DGPs based on richly specified dynamic factor models: VAR confidence intervals fail to adequately control coverage in a sizable fraction of DGPs, while LP confidence intervals robustly maintain accurate coverage. They document that LP is not only more robust to lag length selection, but also to the choice of control variables, consistent with the theory in this paper.
@@ 6 Conclusion
    When the goal is to construct confidence intervals for impulse responses that have accurate coverage in a wide range of empirically relevant DGPs—as opposed to minimizing MSE—then the smaller bias of LPs documented in simulations by Li et al. (2024) is more valuable than the smaller variance enjoyed by VAR estimators.
@@
Is there a way forward for VAR inference, beyond just including a large number of lags? We showed how to construct a VAR confidence interval with a bias-aware critical value that robustly controls coverage, but found that it will typically lead to wider confidence intervals than LP. Another option would be to estimate VARMA models rather than pure VARs, though this would be computationally expensive, and the bias-variance trade-off relative to LPs is unclear. In principle, VAR procedures may work better under additional restrictions on the misspecification, such as shape restrictions on the impulse response functions. [^11] Given any convex parameter space for the misspecification MA polynomial $\alpha(L)$, the worst-case bias of the VAR estimator (see Proposition 3.2) can be computed using convex programming. However, it appears that detailed application-specific restrictions would be required to generate a negligible worst-case bias, since we have shown that the least favorable misspecification in our baseline analysis cannot generally be ruled out based on economic theory. Rather than restricting the parameter space, future research could instead investigate weakening the coverage requirement, e.g., only requiring a certain coverage probability on average over a set of horizons (Armstrong et al., 2022), or by changing the target for inference from the true impulse response function to a smooth projection of this function (Genovese and Wasserman, 2008). Finally, a subjectivist Bayesian VAR modeler need only worry about our negative results if their prior on potential misspecification attaches significant weight to MA processes that imply large VAR biases.